1. The specific heat of aluminium is 0.902 J/g.oC. How much heat is absorbed by an aluminium pie tin with a mass of 473 g to raise its temperature from room temperature (23.00oC) to oven temperature (375oF)?
1
Expert's answer
2020-10-20T08:52:13-0400

T1 = 23.0 ºC

T2 = 375 ºF = 190.55 ºC

∆T = 190.55 – 23.0 = 167.55

m = 473 g

C = 0.902 J/g ºC

∆H = mC∆T

"∆H =473 \times 0.902 \times 167.55 = 71484 \;J = 71.5\; kJ"

Answer: 71.5 kJ